# Optimization of Module Transferability in Single Image Super-Resolution: Universality Assessment and Cycle Residual Blocks

### v1

@@ Abstract
Deep learning has substantially advanced the Single Image Super-Resolution (SISR). However, existing researches have predominantly focused on raw performance gains, with little attention paid to quantifying the transferability of architectural components. In this paper, we introduce the concept of “Universality” and its associated definitions which extend the traditional notion of “Generalization” to encompass the modules’ ease of transferability, thus revealing the relationships between module universality and model generalizability. Then we propose the Universality Assessment Equation (UAE), a metric for quantifying how readily a given module could be transplanted across models. Guided by the UAE results of standard residual blocks and other plug-and-play modules, we further design two optimized modules, Cycle Residual Block (CRB) and Depth-Wise Cycle Residual Block (DCRB). Through comprehensive experiments on natural-scene benchmarks, remote-sensing datasets, extreme-industrial imagery and on-device deployments, we demonstrate that networks embedded with the proposed plug-and-play modules outperform several state-of-the-arts, reaching a PSNR enhancement of up to 0.83dB or enabling a 71.3% reduction in parameters with negligible loss in reconstruction fidelity.
Keywords Generalization; Optimization; Super-resolution; Universality
@@ 1 Introduction
To enable a quantitative assessment of module transferability and provide a novel perspective on plug-and-play module optimization, this paper proposes the definition of “Universality” and its associated concepts. Subsequently, the Universality Assessment Equation is designed by analyzing the pivotal characteristics of existing plug-and-play modules and other state-of-the-art networks. Cycle Residual Blocks (CRB) and Depth-wise Cycle Resiual Blocks (DCRB) are further derived on the basis of RB under the guidance of UAE results. Experiments demonstrate that models embedded with the optimized modules could achieve a 0.83dB increase in PSNR, or reduce parameters by up to 71.3% with minimal performance loss. The main contributions of this paper are as follows:
@@
3) The effectiveness of CRB and DCRB is validated through experiments on various SISR tasks ranging from conventional settings to extreme industrial scenarios (e.g., coalmine images), where the optimized modules outperform several state-of-the-arts and significantly enhance the visual quality of reconstructed images. Moreover, through on-device deployments, we have validated the practicability of CRB in real-world applications.
@@ 2.1 Models for Single Image Super-Resolution
Existing SISR networks could be broadly classified into four categories based on their backbone models: CNN-based models ([1, 6, 4]), GNN-based models ([17, 20, 21, 22]), GAN-based models ([23]), and Transformer-based models ([15, 13, 42, 24]). While highly integrated models could reduce computational costs and improve the performance, they also make it difficult to achieve efficient plug-and-play compatibility across different networks. For instance, the RSTB module in SwinIR [13] contains STL layers, which in turn incorporate MSA (Multi-Head Self-Attention) and MLP (Multi-Layer Perceptron). The deep integration of these components requires careful structure modifications when transferring the module to other backbones.
@@ 2.2 Plug-and-Play Modules in SISR Networks
Existing SISR methods often incorporate various modules into the networks to enhance model performance. For instance, RB, introduced in ResNet [18], has been widely adopted in diverse architectures. ConvFFN, introduced in ViT [16], is used in the GNN-based IPG [17]. SRFlow-DA [27] embeds RB directly into the affine injector layer to improve the performance of SRFlow [28]. These plug-and-play modules significantly enhance model performance but are relatively scarce and difficult to design, since structures of the plug-and-play modules are required to be highly streamlined while preserving an ideal performance. Therefore, we attempt to analyze the structure of plug-and-play modules from a universality perspective, thus designing modules with enhanced performance.
@@ 2.3 Generalization and Universality
“Generalization” denotes a model’s ability to maintain high performance on test sets distinct from its training data [29]. IPT [15], for example, utilizes large-scale pre-training to unify three different low-level vision tasks. DGUNet [19], based on DNN, designs an interpretable SISR network featuring universality. While these studies enhance model performance by improving generalizability, they lack a quantitive evaluation of local module transferability. Therefore, this paper introduces the concept of universality to characterize module transferability, thereby facilitating the design of plug-and-play modules that enhance model generalization.
@@ 3 Proposed Methods
In this section, definitions for “Universality” and its associated concepts are proposed to lay the foundation for a qualitative description of module transferability, namely the Universality Assessment Equation (UAE). We further optimize RB into CRB/DCRB through UAE analysis and the optimization mechanism under UAE guidance is elucidated from the perspective of back propagation. Finally, a generalized designing principle for CRB is explained to provide a clearer picture on module optimization.
@@
Universality Local module Complete model The ease of module transferability
Generalization Complete model Brand new test data Whether model performance remains ideal
@@
Universality and generalization are not completely isolated. Differences between the previously proposed plug-and-play modules ([18, 11, 31, 32]) reveal that while more complex blocks could achieve superior generalizability, they also typically incur larger parameter counts and more elaborate computational graphs, all of which may deteriorate the ease of transferability. In other words, model’s generalizability could be enhanced by appropriately decreasing the module universality.
@@ 3.2 Universality Assessment Equation (UAE)
Figure 2: Structures of the proposed CRB and DCRB, where $l$ stands for the UAE variable.
@@ 3.3 Cycle Residual Block and Depth-wise Separable Cycle Residual Block
Let $X\in{R}^{B\times C\times H\times W}$ denote the input tensor, where $B$, $C$, $H$ and $W$ correspond to batch size, channel dimensions, height and width, respectively. The CRB and its depth-wise variant DCRB are then defined as Eq.(3), where Conv($\cdot$) represents standard convolution for CRB and depth-wise separable convolution for DCRB.
@@
where $L(x)$ stands for losses and $J_{i}$ the Jacobian matrix. Let $\mathbf{F}_{l}(\cdot):\mathbb{R}^{d}\xrightarrow{}\mathbb{R}^{d}$ denote the nonlinear mapping (e.g., convolution followed by activation ) in the $l$-th residual block. Since each block implements $x_{l+1}=x_{l}+\mathbf{F}_{l}(x_{l})$, the total derivative of the loss with respect to initial input $x_{0}$ follows by the repeated application of chain rule:
@@
Figure 3: Experimental validation of Eqs.(5), (7) and (12), where EDSR [6] is chosen as the backbone. EDSR$+$CRB attains 31.00 dB PSNR by epoch 200, whereas EDSR$+$RB achieves the same value around epoch 350, confirming CRB’s faster convergence.
Figure 4: Experimental validation of Eqs.(14) and (16), where EDSR [6] is chosen as the backbone. CRBl=16,ε=4 exhibits smooth convergence, whereas CRBl=16,ε=5 fails to converge and possesses a substantially degraded performance, with its loss plateauing at approximately 632.5 during the experiments.
Let $A,B\in\mathbb{R}^{d\times d}$ satisfy: (1) “$A$ and $B$ are symmetric positive semidefinite matrices”, (2) “$A$ and $B$ commute, i.e. $AB=BA$”. Thus, there exists orthogonal matrix $Q$ satisfying Eq.(9).
@@
[…]
@@
Such approximation is grounded in existing physical principles. From the perspective of Neural-ODEs [33], a residual block implements one step of the Euler method: $x(t+\Delta t)=x(t)+\Delta t\cdot f(x(t))$, yielding $F(x)=x(t+\Delta t)-x(t)\approx\Delta t\cdot f(x)$ and $F^{\prime}\xrightarrow{}0$. Treating $F^{\prime}$ as a “small perturbation” aligns with the fact that residual learning ultimately guides deep modules degenerate into identity function. In quantum mechanics, Dysen series and Lie-Trotter product formulas [34] use the approximation $e^{\frac{A}{n}}\approx I+\frac{1}{n}A+O(n^{-2})$ retaining merely the linear “$I+$small” at each step, motivating the requirements $\|E+A\|_{2}\geq 1$ in Eq.(12).
Therefore, even relaxing the symmetric/commuting conditions, Eq.(12) together with Eq.(7) indicate that every added residual branch contributes at least a unit of spectral gain1absent1$\geq$ 1≥ 1 into the back propagation, leading to smoother gradient flow, faster convergence, and better module performance, as is verified through experiments in Fig.3.
@@
Table 3: Quantitative comparison results (PSNR/SSIM) between several universal modules with 4×4$\times$4 × SR scale, trained on DIV2K. Bold texts indicate the best performance.
@@
[…]
@@ 3.4 Generalized Forms for CRBs and its Designing Principles
Although Eq.(12) demonstrates that introducing extra residual connections accelerates module convergence without adding extra parameters, we derive an upper bound of $\frac{l}{4}$ on the residual number $\varepsilon$ ($\varepsilon\leqslant\frac{l}{4}$), where $l$ is the CRB’s UAE variable and the constant 4 equals half the number of forward propagation layers per CRB.
Following Eq.(7), we initially verify that for $\varepsilon=4,l=16$, CRB converges stably. From Fig.2, CRB ($\varepsilon=4,l=16$) could be expressed as Eq.(13)
$y_{1}=x+\mathbf{F}_{1}(x),y_{2}=y_{1}+\mathbf{F}_{2}(y_{1}),y_{3}=y_{2}+\mathbf{F}_{3}(y_{2}),y_{4}=y_{3}+\mathbf{F}_{4}(y_{3}).$ (13)
Hence, for a network comprised of $L$ CRBε=4, the back propagation is given by Eq.(14). Fig.LABEL:converge_1 and Fig.LABEL:loss_1 provides the experimental validation for the convergence of Eq.(14).
$\frac{\partial L}{\partial x_{0}}=\frac{\partial L}{\partial x_{L}}\prod_{l=0}^{L-1}\Big{[}(\mathbf{E}+\frac{\partial\mathbf{F}_{4,l}}{\partial y_{3,l}})(\mathbf{E}+\frac{\partial\mathbf{F}_{3,l}}{\partial y_{2,l}})(\mathbf{E}+\frac{\partial\mathbf{F}_{2,l}}{\partial y_{1,l}})(\mathbf{E}+\frac{\partial\mathbf{F}_{1,l}}{\partial x_{l}})\Big{]}=\frac{\partial L}{\partial x_{L}}\prod_{i=0}^{L-1}J_{l}^{\text{CRB}_{\varepsilon=4}}\end{aligned}.$ (14)
However, when an additional residual connection is introduced to Eq.(13), as shown in Eq.(15), the back propagation should be updated as Eq.(16).
$y_{1}=y_{0}+\mathbf{F}_{1}(y_{0}),y_{2}=y_{1}+\mathbf{F}_{2}(y_{1})+x,y_{3}=y_{2}+\mathbf{F}_{3}(y_{2}),y_{4}=y_{3}+\mathbf{F}_{4}(y_{3}).$ (15)
$\frac{\partial L}{\partial x_{0}}=\frac{\partial L}{\partial x_{L}}\Bigg{[}\prod_{l=0}^{L-1}\Big{[}\prod_{i=1}^{4}(\mathbf{E}+\frac{\partial\mathbf{F}_{i,l}}{\partial y_{i-1,l}})+\Theta_{l}\Big{]}\Bigg{]}=\frac{\partial L}{\partial x_{L}}\prod_{l=0}^{L-1}J_{l}^{\text{CRB}_{\varepsilon=5}},\Theta_{l}=\prod_{t=3}^{4}(\mathbf{E}+\frac{\partial\mathbf{F}_{t,l}}{\partial y_{t-1,l}}).$ (16)
In conjunction with Eq.(12), it is clear that $\|J_{l}^{\text{CRB}_{\varepsilon=5}}\|_{2}>2>\|J_{l}^{\text{CRB}_{\varepsilon=4}}\|_{2}>1$, indicating that $\text{CRB}_{l=16,\varepsilon=5}$ is highly susceptible to trigger gradient exploding. Such conclusion derived from theoretical analysis are corroborated by the experimental results presented in Fig.LABEL:converge_2 and Fig.LABEL:loss_2, where $\text{CRB}_{l=16,\varepsilon=5}$ exhibits gradient explosions in first 20 epochs accompanied by a substantial decline in PSNR performance.
We now generalize the setting to a CRB with $l$ forward propagation layers where $\varepsilon$ residual connections are introduced, as shown in Eq.(17). Assuming that the previously stated hypothesis on the upper bound of residual numbers $\varepsilon\leqslant\frac{l}{4}(\varepsilon\in\mathbb{N})$ holds, then the computed results should theoretically indicate that the module converges for $\varepsilon=\frac{l}{4}$ but fails to converge for $\varepsilon=\frac{l}{4}+1$.
$\begin{cases}y_{1}=x+\mathbf{F}_{1},\cdots,\ y_{\varepsilon}=y_{\varepsilon-1}+\mathbf{F}_{\varepsilon},&\varepsilon=\frac{l}{4}\\ y_{1}=x+\mathbf{F}_{1},\ y_{2}=y_{1}+\mathbf{F}_{2}+x,\cdots&\varepsilon=\frac{l}{4}+1\end{cases}$ (17)
Referring to Eqs.(14) and (16), we derive the back propagation process ($\varepsilon=\frac{l}{4}$ for Eq.(18) and $\varepsilon=\frac{l}{4}+1$ for Eq.(19)) for the model made up of $L$ generalized CRBs. Similar to Eq.(12), it could be easily concluded that $\|J_{l}^{\text{CRB}_{\varepsilon=l/4+1}}\|_{2}>2>\|J_{l}^{\text{CRB}_{\varepsilon=l/4}}\|_{2}>1$. Consequently, when $\varepsilon>\frac{l}{4}$, $\prod_{l=0}^{L-1}J_{l}^{\text{CRB}_{\varepsilon}}$ exhibits exponential growth with a base larger than 2, leading to gradient explosion and declined module performance.
$\frac{\partial L}{\partial x_{0}}=\frac{\partial L}{\partial x_{L}}\prod_{l=0}^{L}\prod_{i=1}^{\varepsilon}(\mathbf{E}+\frac{\partial\mathbf{F}_{i,l}}{\partial y_{i-1,l}})=\frac{\partial L}{\partial x_{L}}\prod_{l=0}^{L-1}J_{l}^{\text{CRB}_{\varepsilon}}.$ (18)
$\frac{\partial L}{\partial x_{0}}=\frac{\partial L}{\partial x_{L}}\Bigg{[}\prod\limits_{l=0}^{L-1}\Big{[}\prod\limits_{i=1}^{\varepsilon}(\mathbf{E}+\frac{\partial\mathbf{F}_{i,l}}{\partial y_{i-1,l}})+\Theta_{l}\Big{]}\Bigg{]}=\frac{\partial L}{\partial x_{L}}\prod_{l=0}^{L-1}J_{l}^{\text{CRB}_{\varepsilon}},\Theta_{l}=\prod_{t=3}^{\varepsilon}(\mathbf{E}+\frac{\partial\mathbf{F}_{t,l}}{\partial y_{t-1,l}}).$ (19)
@@
Table 5: Quantitative results on Potsdam remote sensing dataset [43] under ×$\times 4$ SR scale. Hyper-parameter settings of IPG and SRFormer are scaled down to accommodate limited computational resources and accelerate training, while EDSR and SRResnet–both featuring moderate parameter counts–retain their full hyper-parameter configurations. The first 700 images of the dataset are used for model training, while the remaining 300 images for performance evaluation.
@@
Table 6: Quantitative results on BelT coalmine dataset [44] under ×$\times 4$ SR scale. Bold texts indicate the best performance.
Models PSNR(dB) SSIM LPIPS
EDSR (CVPR 2017) [6] 28.06 0.7551 0.1154
SRFormer (ICCV 2023) [42] 29.76 0.7906 0.1017
IPG (CVPR 2024) [17] 29.95 0.7921 0.1028
EDSR (with CRB) 28.54 0.7604 0.1122
IPG (with CRB) 30.03 0.7930 0.1010
IPG (with DCRB) 29.97 0.7928 0.0999
SRFormer (with CRB) 29.81 0.7884 0.1042
SRFormer (with DCRB) 29.80 0.7913 0.1033
In summary, the generalized CRB converges stably and achieves optimal performance only when $\varepsilon\leqslant\frac{l}{4}$ is satisfied, where $\varepsilon$ is the number of residual connections and $l$ the number of UAE propagation layers (Eq.(1)).

### v2

@@ Abstract
Deep learning has substantially advanced the Single Image Super-Resolution (SISR). However, existing researches have predominantly focused on raw performance gains, with little attention paid to quantifying the transferability of architectural components. In this paper, we introduce the concept of “Universality” and its associated definitions which extend the traditional notion of “Generalization” to encompass the modules’ ease of transferability. Then we propose the Universality Assessment Equation (UAE), a metric which quantifies how readily a given module could be transplanted across models and reveals the combined influence of multiple existing metrics on transferability. Guided by the UAE results of standard residual blocks and other plug-and-play modules, we further design two optimized modules, Cycle Residual Block (CRB) and Depth-Wise Cycle Residual Block (DCRB). Through comprehensive experiments on natural-scene benchmarks, remote-sensing datasets and other low-level tasks, we demonstrate that networks embedded with the proposed plug-and-play modules outperform several state-of-the-arts, reaching a PSNR enhancement of up to 0.83dB or enabling a 71.3% reduction in parameters with negligible loss in reconstruction fidelity. Similar optimization approaches could be applied to a broader range of basic modules, offering a new paradigm for the design of plug-and-play modules.
Keywords Generalization; Optimization; Super-resolution; Universality; Plug-and-Play
@@ 1 Introduction
To enable a quantitative assessment of module transferability and provide a novel perspective on plug-and-play module optimization, this paper proposes the definition of “Universality” and its associated concepts. Subsequently, the Universality Assessment Equation (UAE) is designed by analyzing the pivotal characteristics of existing plug-and-play modules and other state-of-the-art networks. Cycle Residual Blocks (CRB) and Depth-wise Cycle Residual Blocks (DCRB) are further derived on the basis of RB under the guidance of UAE results. Experiments demonstrate that models embedded with the optimized modules could achieve a 0.83dB increase in PSNR, or reduce parameters by up to 71.3% with minimal performance loss. The main contributions of this paper are as follows:
@@
3) The effectiveness of CRB and DCRB is validated through experiments on various SISR tasks, and the UAE optimization strategy is further verified through evaluation on various low-level tasks such as denoising and deblurring. Experiments demonstrate that our optimization strategy features strong generalizability across different vision tasks, offering a new paradigm for the optimization of plug-and-play modules.
@@ 2.1 Models for Single Image Super-resolution
Early SISR networks could be broadly classified into four categories based on their backbone models: CNN-based models ([1, 6, 4]), GNN-based models ([22, 25, 26, 27]), GAN-based models ([28]), and Transformer-based models ([20, 13, 46, 29]). While highly integrated models could reduce computational costs and improve the performance, they also make it difficult to achieve efficient plug-and-play compatibility across different networks. For instance, the RSTB module in SwinIR [13] contains STL layers, which in turn incorporate MSA (Multi-Head Self-Attention) and MLP (Multi-Layer Perceptron).
Recent hybrid architectures introduce new paradigms but compromise transferability as well. SRMamba-T [14] merges Mamba’s sequential scanning with Transformer’s global attention for long-range dependency modeling. However, the fundamental difference between Mamba’s linear scanning mechanism and Transformer’s global self-attention necessitates intricate signal pathways, resulting in considerable architectural coupling. HSR-KAN [15] integrates KANs (Kolmogorov-Arnold Networks) with CNN and MLP to reach a trade-off between efficiency and quality , but its specialized fusion structure (e.g., KAN-CAB module) severely limits its task generalizability. Although the BUFF (Bayesian Uncertainty Guided Diffusion Model) [16] introduces Bayesian methods into diffusion models to relax the assumption of independent noises, however, the instability of Bayesian training substantially increases model coupling and computational intensity. Furthermore, the tight coupling in diffusion models between the forward diffusion and reverse denoising processes precludes any individual component from functioning independently. Therefore, while these advances boost performance, their intricate integrations impede module reuse across architectures.
@@ 2.2 Plug-and-Play Modules in SISR Networks
The concept of plug-and-play modules emerges from the recognition that modular components could improve reproducibility across different SISR architectures. Early successful examples include the widespread use of RB across diverse networks, from CNN-based EDSR [6] to GAN-based ESRGAN [28], demonstrating the value of structurally simple yet effective components.
The attention mechanisms represent one of the most successful transferable modules in recent years. Beyond RCAB’s [11] channel attention, researchers developed increasingly sophisticated attention variants. For instance, MDAB (Multi-scale Dilated Attention Block) [17] offers a lightweight attention mechanism. However, its effectiveness critically depends on the pre-processing by a dedicated LRM (Local Residual Module). Consequently, transplanting MDAB without its coupled LRM generally leads to performance degradation.
These plug-and-play modules significantly enhance model performance but remain relatively scarce and challenging to design, since they demand highly streamlined structures while preserving its ideal performance. Therefore, analyzing the module structures from a universal perspective is crucial for designing improved plug-and-play modules.
@@ 2.3 Generalization and Universality
“Generalization” refers to a model’s capacity to maintain high performance when evaluated on data distinct from its training distribution [34]. Existing researches exploring “universality” predominantly focus on the developments of unified frameworks and task formulations. For instance, IPT [20] leverages large-scale pre-training to unify diverse low-level vision tasks, while DGUNet [24] proposes an interpretable, unified deep neural network architecture. Similarly, Uni-COAL[18] introduces a unified framework for Magnetic Resonance Imaging super-resolution. Although these approaches enhance model performance by improving framework universality, they have not thoroughly investigated or provided quantitative evaluation for the module transferability. Therefore, this paper proposes the concept of universality to characterize the transferability of individual components. This quantification facilitates the systematic design of plug-and-play modules aimed at enhancing model generalization and framework universality.
@@ 3 Proposed Methods
In this section, definitions for “Universality” and its associated concepts are proposed to lay the foundation for a quantitative description of module transferability, namely the Universality Assessment Equation. We further optimize RB into CRB/DCRB through UAE analysis and the optimization mechanism under UAE guidance is elucidated from the perspective of back propagation. Finally, a generalized designing principle for CRB is explained to provide a clearer picture on module optimization.
@@
Universality
Local
module
Complete
model
The ease of
module transferability
Generalization
Complete
model
Brand new
test data
Whether model
performance remains ideal
@@
Universality and generalization are not completely isolated. Differences between the previously proposed plug-and-play modules ([23, 11, 36, 37]) reveal that while more complex blocks could achieve superior generalizability, they also typically incur larger parameter counts and more elaborate computational graphs, all of which may deteriorate the ease of transferability. Therefore, generalizability could be enhanced by appropriately decreasing the module universality.
@@ 3.2 Universality Assessment Equation (UAE)
Figure 2: Overview of the proposed blocks and the UAE optimization strategy.
@@ 3.3 Cycle Residual Block and Depth-wise Separable Cycle Residual Block
Let $X\in\mathbb{R}^{B\times C\times H\times W}$ denote the input tensor, where $B$, $C$, $H$ and $W$ correspond to batch size, channel dimensions, height and width, respectively. The CRB and its depth-wise variant DCRB are then defined as Eq.(3), where Conv($\cdot$) represents standard convolution for CRB and depth-wise separable convolution for DCRB.
@@
where $L(x)$ stands for losses and $J_{i}$ the Jacobian matrix. Let $\mathbf{F}_{l}(\cdot)$ denote the nonlinear mapping (e.g., convolution followed by activation ) in the $l$-th residual block. Since each block implements $x_{l+1}=x_{l}+\mathbf{F}_{l}(x_{l})$, the total derivative of the loss with respect to initial input $x_{0}$ follows by the repeated application of chain rule:
@@
Let $A,B\in\mathbb{R}^{d\times d}$ satisfy: (1) “$A$ and $B$ are symmetric positive semidefinite matrices”, (2) “$A$ and $B$ commute, i.e. AB=BAAB=BA”. Thus, there exists orthogonal matrix $Q$ satisfying Eq.(9).
@@
$\|E+A\|_{2}$ $=\underset{i}{\max}(1+\alpha_{i}),\|(E+A)(E+B)\|_{2}$ $=\underset{i}{\max}[(1+\alpha_{i})(1+\beta_{i})].$ (11)
where $\underset{i}{\max}\beta_{i}\triangleq\beta_{\max}\geq 0$. Finally we arrive at the expression given in Eq.(12).
@@
Such approximation is grounded in existing physical principles. From the perspective of Neural-ODEs [38], a residual block implements one step of the Euler method: $x(t+\Delta t)=x(t)+\Delta t\cdot f(x(t))$, yielding $F(x)=x(t+\Delta t)-x(t)\approx\Delta t\cdot f(x)$ and $F^{\prime}\xrightarrow{}0$. Treating $F^{\prime}$ as a “small perturbation” aligns with the fact that residual learning ultimately guides deep modules degenerate into identity function.
[…]
@@
Table 3: Quantitative comparison results (PSNR/SSIM) between several universal modules with $4\times$ SR scale, trained on DIV2K.
@@
Table 4: Quantitative comparison (PSNR/SSIM/LPIPS) for RB [6] and CRB/DCRB, using EDSR [6] as the backbone. Oursii represents the incorporation of corresponding module into the baseline.
@@ 3.4 Generalized Forms for CRBs and its Designing Principles
Although Eq.(12) demonstrates that introducing extra residual connections accelerates module convergence without adding extra parameters and is beneficial for performance improvements, we derive an upper bound on the residual number $\varepsilon$ ($\varepsilon\leqslant l/4+\sum\varepsilon_{x}\leqslant l/4+\delta_{\text{safe}},\delta_{\text{safe}}\geq 0$), where $l$ is the UAE variable inferred from computational graphs, $\delta_{\text{safe}}$ is correlated with hyperparameters and experimental settings, and $\varepsilon_{x}$ represents the extra residual connections introduced by inputs. Additionally, we theoretically demonstrate that $\sum\varepsilon_{x}$ is inversely proportional to the module’s stability. When $\sum\varepsilon_{x}>\delta_{\text{safe}}$, the module undergoes gradient explosion. In our ablation study, we reveal the nature of this phenomenon and clarify why the module performance degrades.
Consider a CRB* as illustrated in Eq.(13), whose $l/4=4,\sum\varepsilon_{x}=|m|+|n|+|z|$. Similar to Eq.(7), we derive the back propagation of a model made up of $n$ CRB*s, as shown in Eq.(14).
$y_{1}=x+\mathbf{F}_{1},y_{2}=y_{1}+\mathbf{F}_{2}+mx,y_{3}=y_{2}+\mathbf{F}_{3}+nx,y_{4}=y_{3}+\mathbf{F}_{4}+zx$ (13)
$\frac{\partial L(x)}{\partial x_{0}}$ $=\frac{\partial L(x)}{\partial x_{n}}\prod_{l=0}^{n-1}\Big{[}\prod_{i=1}^{4}(\mathbf{E}+\frac{\partial\mathbf{F}_{i,l}}{\partial y_{i-1,l}})+m\prod_{i=3}^{4}(\mathbf{E}+\frac{\partial\mathbf{F}_{i,l}}{\partial y_{i-1,l}})+n(\mathbf{E}+\frac{\partial\mathbf{F}_{4,l}}{\partial y_{3,l}})+z\mathbf{E}\Big{]}$ (14)
$=\frac{\partial L(x)}{\partial x_{n}}\prod_{l=0}^{n-1}J_{l}^{\text{CRB*}}$
It is obvious that $1+\sum\varepsilon_{x}<\|J_{l}^{\text{CRB*}}\|_{2}<2+\sum\varepsilon_{x}$, where $\sum\varepsilon_{x}=|m|+|n|+|z|$. To reveal the threshold conditions that prevent modules from experiencing gradient explosion, we consider an extreme condition where $\sum\varepsilon_{x}=\delta_{\text{safe}}$, leading to: $\|\partial L(x)/\partial x_{0}\|_{2}\in\Big{(}(1+\delta_{\text{safe}})^{n},(2+\delta_{\text{safe}})^{n}\Big{)}$. Therefore, to minimize the risk of gradient explosion, it is necessary to set $\delta_{\text{safe}}=0$. When $\delta_{\text{safe}}>0$, various factors (e.g., hyperparameters) collectively govern the module’s convergence, and the risk of gradient explosion escalates as $\delta_{\text{safe}}$ increases.
To generalize our conclusion, we further consider a CRB with $l$ forward propagation layers where $\varepsilon$ residual connections are introduced, as shown in Eq.(15).
$y_{1}=x+\mathbf{F}_{1},y_{2}=y_{1}+\mathbf{F}_{2}+c_{1}x,y_{3}=y_{2}+\mathbf{F}_{3}+c_{2}x,\cdots,\ y_{\varepsilon}=y_{\varepsilon-1}+\mathbf{F}_{\varepsilon}+c_{\varepsilon-1}x$ (15)
Referring to Eqs.(13) and (14), we derive the back propagation process for the model made up of $L$ generalized CRBs.
By defining $D_{k}=\partial y_{k}/\partial x$, the recurrence relation for $D_{k}$ is established as Eq.(16).
$D_{k}=(\mathbf{E}+\frac{\partial\mathbf{F}_{k}}{\partial y_{k-1}})D_{k-1}+c_{k-1}\mathbf{E},k\geq 2$ (16)
By calculating Eq.(16) , the general form of $D_{k}$ goes as:
$D_{k}=\prod_{i=1}^{k}(\mathbf{E}+\frac{\partial\mathbf{F}_{i}}{\partial y_{i-1}})+\sum_{j=1}^{k-1}c_{j}\prod_{i=j+2}^{k}(\mathbf{E}+\frac{\partial\mathbf{F}_{i}}{\partial y_{i-1}})+c_{k-1}\mathbf{E}$ (17)
When setting $k=l$, we derive the Jacobian matrix of generalized CRB:
$J_{l}=D_{k=l}=\prod_{i=1}^{l}(\mathbf{E}+\frac{\partial\mathbf{F}_{i}}{\partial y_{i-1}})+\sum_{j=1}^{l-1}c_{j}\prod_{i=j+2}^{l}(\mathbf{E}+\frac{\partial\mathbf{F}_{i}}{\partial y_{i-1}})+c_{l-1}\mathbf{E}$ (18)
Similarly, $1+\sum_{j=1}^{l-1}|c_{j}|=1+\sum\varepsilon_{x}<\|J\|_{2}<2+\sum\varepsilon_{x}$. Consequently, to most straightforwardly ensure that CRB could enhance performance with the least probability of encountering gradient explosion, $\delta_{\text{safe}}$ can be simply set to zero.
@@
Table 5: Quantitative results on Potsdam remote sensing dataset [47] under $\times 4$ SR scale. Hyper-parameter settings of IPG and SRFormer are scaled down to accommodate limited computational resources and accelerate training, while EDSR and SRResnet–both featuring moderate parameter counts–retain their full hyper-parameter configurations. The first 700 images of the dataset are used for model training, while the remaining 300 images for performance evaluation.